A MOLECULAR BEAM OPTICAL STARK STUDY OF THE [15.8] AND [16.0] ^{2}Π_{1/2} − X^{4}Σ^{−} (0,0) BAND SYSTEMS OF RHODIUM MONOXIDE, RhO.

\underline{JAMIE GENGLER}, TONGMEI MA, TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; ALLAN G. ADAM, Department of Chemistry, University of New Brunswick, Fredericton, NB E3B 6E2.

A supersonic molecular beam sample of rhodium monoxide, RhO, has been generated using a laser ablation/chemical reaction generation scheme and probed using near natural line width limited optical Stark spectroscopy utilizing laser induced fluorescence detection. Stark splittings in the ^{S}R_{11}(0) and ^{S}R_{11}(1) lines of the [15.8] and [16.0] ^{2}Π_{1/2} − X^{4}Σ^{−} (0,0) transitions were observed upon application of static electric fields up to 2400 V/cm. Comparisons of the determined field-free molecular parameters is made with previous work^{a}’^{b}’^{c}. Progress on the analysis of the Stark effect will be reported.

aR.F. Heuff, W.J. Balfour, and A.G. Adam, J. Mol. Spectrosc. 216, 136-150 (2002).

bR.H. Jensen, S.G. Fougère, and W.J. Balfour, Chem. Phys. Lett. 370, 106-111 (2003).

cR.F. Heuff, S.G. Fougère, and W.J. Balfour, J. Mol. Spectrosc. 231, 99-107 (2005).